# TESS discovery of two super-Earths orbiting the M-dwarf stars TOI-6002 and TOI-5713 near the radius valley

### v2

@@
We present the validation of two TESS super-Earth candidates transiting the mid-M dwarfs TOI-6002 and TOI-5713 every 10.90 and 10.44 days, respectively. The first star (TOI-6002) is located pc away, with a radius of , a mass of and an effective temperature of K. The second star (TOI-5713) is located pc away, with a radius of , a mass of and an effective temperature of K. We validated the planets using TESS data, ground-based multi-wavelength photometry from many ground-based facilities, as well as high-resolution AO observations from Keck/NIRC2. TOI-6002 b has a radius of and receives . TOI-5713 b has a radius of and receives . Both planets are located near the radius valley and near the inner edge of the habitable zone of their host stars, which makes them intriguing targets for future studies to understand the formation and evolution of small planets around M-dwarf stars.
@@
The abundance and the distribution of planets with radii below 4 for orbits shorter than 100 days allows the identification of specific features that can link their abundance to their formation and evolution like the radius valley (see e.g. Fulton et al., 2017; Cloutier & Menou, 2020) and the density valley (Luque & Pallé, 2022), based on their radii and densities distributions. The radius valley features a lack in planets ranging from 1.5 to 2 . In this regard, some studies argue that rocky super-Earths are a byproduct of atmospheric escape experienced by gaseous sub-Neptunes driven by the energy radiated by the host stars (see e.g. Lopez & Fortney, 2013; Owen & Wu, 2013; Jin et al., 2014; Chen & Rogers, 2016) and/or the energy released by their cooling planetary cores (see e.g. Owen & Wu, 2013; Ginzburg et al., 2016, 2018). Other studies argue that some planets are formed as rocky super-Earths in a gas-depleted (see e.g. Lopez & Rice, 2018; Cloutier & Menou, 2020; Cherubim et al., 2023) or gas-poor environment with thin gaseous envelopes (see e.g. Lee & Connors, 2021; Cherubim et al., 2023). Additional theoretical studies predict that the sub-population of sub-Neptunes consists typically of water-rich worlds formed beyond the water ice-line and migrate toward their host stars (se e.g. Venturini et al., 2020; Burn et al., 2024; Venturini et al., 2024). The density valley features three peaks in the density distribution defining rocky super-Earths, water-rich worlds, and puffy sub-Neptunes. Luque & Pallé (2022) suggest that the pebble accretion is the main mechanism forming these planets with super-Earths formed within the ice line and water-rich world formed outside it (se e.g. Venturini et al., 2020; Burn et al., 2024; Venturini et al., 2024). For now the mechanisms forming super-Earths remain unclear, making observations of planets in this mass-radius regime a critical part to understand it.
TESS is increasing the number of small exoplanets amenable to mass and atmospheric characterization (Ricker et al., 2015; Sullivan et al., 2015). Detailed studies of such exoplanets will allow insights into the origin of the radius/density valley in the near future. Such analyses should shed further light on the mechanisms dominating the formation and evolution of super-Earths. In addition, the analyses of exoplanets with incident irradiation between Earth and Venus will allow key insights into the limits of the HZ (e.g. Kaltenegger et al., 2023a).
@@
The paper is structured as follows. Sect. 2 introduces the data from TESS and all ground-based observations. Stellar characterization, validation of the planetary nature of the two TESS candidates, and global data analyses are covered in Sect. 3. Results and potential for mass measurements and atmospheric characterization are discussed for both planets in Sect. 4. Sect. 5 summarizes our findings.
@@ 2.1 TESS photometry
The TESS time-series observations were processed by the TESS Science Processing Operations Center (SPOC) pipeline (Jenkins et al., 2016, 2020). SPOC calibrates the image data, carries out quality control, labels bad data, and performs simple aperture photometry (SAP) for each star in the field of view. Then, it corrects for instrumental systematic offsets using a Presearch Data Conditioning module (PDC: Smith et al., 2012; Stumpe et al., 2012, 2014) and produces pre-search data conditioned light curves. Fig. 1 shows the target pixel files (TPFs) and the SAP apertures used for photometry, along with the overlaid locations of nearby Gaia DR2 (Gaia Collaboration et al., 2018) sources.
@@
For our analyses, we downloaded the 2-minute PDC-SAP light curves from the Mikulski Archive for Space Telescopes (MAST111https://archive.stsci.edu/). Using the ”QUALITY” flag, we identified all the bad-quality data points that were removed from the data files. The year two data of TESS were subject to sky background overestimation in the SPOC pipeline for dim and/or crowded targets, affecting the transit depth and, thus, the planet radius. TOI-6002 is a crowded-field target and was subjected to this effect in sector 14. We corrected the PDC-SAP light curves of sector 14 using equation 2 described in section 4.2 of the data release notes of sector 27222https://tasoc.dk/docs/release_notes/tess_sector_27_drn38_v02.pdf (see section 4.2, equation 2). This effect was negligible for TOI-5713 due to the low level of crowding. The light curve of TOI-5713 showed some clear flare-like signals, which we removed using an upper 3 clipping. We used wotn (Hippke et al., 2019) to detrend the light curves for stellar variability, utilizing a bi-weight filter with a window 2.5 times longer than each planets’ transit duration. The obtained lightcurves are presented in Fig. 2, with the locations of the transit signals highlighted with different colors.
@@ 2.2.1 LCOGT-1m0 photometry
We used the 40964096 LCOGT SINISTRO CCD camera on the 1.0-m robotic telescope of the Las Cumbres Observatory Global Telescope (LCOGT; Brown et al. 2013) network node at Teide Observatory (Tenerife, Spain) to observe two full transits of TOI-6002 b on UT 2023 May 13 and 24. Table 1 provides the observation parameters. We used the standard LCOGT BANZAI pipeline (McCully et al., 2018) to calibrate the science images, and we used AstroImageJ (AIJ: Collins et al., 2017) for the photometric extraction.
@@ 2.2.4 MuSCAT2 photometry
For TOI-5713 b, we observed one full transit on UTC July 8, 2023 (see Table 1), using the multi-color imager (, , , and bands) MuSCAT2 (Narita et al., 2019), mounted on the 1.52-m Telescopio Carlos Sánchez (TCS) at Teide Observatory, Spain. We used the band images for guiding, as the star is faint in the blue optical ( mag). We calibrated and reduced the raw data using the code and procedure of (Parviainen et al., 2019). Due to the short exposure time and the faintness of the target, the MuSCAT2 band photometry presented a much larger scatter than in the other bands, and it was not used in the analysis presented below.
@@ 2.3.1 IRTF/SpeX
We used the SpeX spectrograph (Rayner et al., 2003) on the 3.2-m NASA Infrared Telescope Facility (IRTF) to collect a near-infrared spectrum of TOI-5713 on 2023 May 16 (UT) and TOI-6002 on 2023 Oct 31 (UT). On both nights, conditions were clear with seeing of 06–07. We used the short-wavelength cross-dispersed (SXD) mode, and aligned the slit to the parallactic angle. The resulting spectra span 0.80–2.42 m and have a resolving power of . We collected six integrations of 139.9 s on TOI-5713 and 299.8 s on TOI-6002, using an ABBA nod pattern. Immediately after the science observations, we gathered a standard set of flat-field and arc-lamp calibrations and observed A0 V standard stars. We reduced the data with Spextool v4.1 (Cushing et al., 2004), following the standard instructions from the Spextool User’s Manual333Available at http://irtfweb.ifa.hawaii.edu/~spex/observer/.. The final spectrum of TOI-5713 has a median per-pixel SNR of 92 with peak values in the , , and bands of of 116, 131, and 121, respectively. For TOI-6002, the final spectrum has a median SNR per pixel of 156 with peak values of 195, 219, and 205 in the , , and bands, respectively.
@@ 2.4 High-Resolution Imaging
We assessed the sensitivities of the final combined AO image by injecting simulated sources azimuthally around the primary target, following the approach of Furlan et al. (2017). Fig 3 shows the resulting sensitivity curves. We do not detect close-in () stellar companions for either star. The Keck and Palomar data to have sensitivities of mag and mag at 0.5″, respectively. This corresponds to a spatial limit au for TOI-6002 and au for TOI-5713.
@@ 3.2 Discarding possible false-positive scenarios
The second astrophysical false positive case is that of bound eclipsing binaries. Many findings argue against this scenario for both stars: (1) the spectral energy distribution is fit well by single-star models (see Fig. 6); (2) the spectra collected by SpeX SXD and Shane/Kast, for both stars, show no secondary spectrum (see Sect. 2), and Fig. 4 and 5; (3) the achromatic nature of the transits, where their depths in all bands agree within for each planet (see Tables 5 and 6); (4) the stellar densities determined from the transits alone are in good agreement with the those determined from the the basic parameters of each star (see Table 3). Considering these arguments, we discard these false-positive scenarios.
@@ 3.6 Global modeling
In the priors file, we set starting values of , and . We set lower and upper limits priors on the orbital period ( d) and the transit epoch () from the values determined by SPOC and reported in ExoFOP. For the LD parameters, we disabled the priors based on the Claret tables (Claret, 2017; Claret & Bloemen, 2011) by setting the NOCLARET flag given the low mass of our stars where these tables are not reliable (see Eastman et al., 2019, and references therin). Instead, we fitted for quadratic LD parameters assuming Gaussian priors of variance 0.1, computed with the PyLDTk code (Parviainen, 2015) for each passband (see Table 4). Regarding the TESS data, we used the detrended PDC-SAP light curves (see Sect. 2.1) which are corrected for the contribution of neighboring stars and subsequently corrected for sky background bias for TOI-6002 in sector 14 (see Sect. 2.1), but we still fit for a possible dilution by applying a Gaussian prior of 010% of the contamination ratio reported in ExoFOP. This is to propagate any uncertainty in the correction. Also, Gaussian priors were applied on the metallicities from IRTF/SpeX (see Table 3). For each planet, we fitted for circular and eccentric orbits to check for orbital eccentricity evidences using only photometric data. The two fits were run until convergence. We found the difference in the loglikelihood is for both TOI-6002 b and TOI-5713 b, which suggests that eccentric orbits are marginally favored for both planets. Tables 5 and 6 present the fits’ results for both the circular and eccentric cases.
@@
| Parameter | Description | Circular | Eccentric |
| Stellar Parameters: | |||
| Mass () | |||
| Radius () | |||
| Luminosity () | |||
| Density (cgs) | |||
| Surface gravity (cgs) | |||
| Effective Temperature (K) | |||
| Planetary Parameters: | |||
| Orbital period (days) | |||
| Planet radius () | |||
| Transit time () | |||
| Semi-major axis (AU) | |||
| Inclination (Degrees) | |||
| Eccentricity | 0 (Fixed) | ||
| Tidal circularization timescale (Gyr) | |||
| Argument of Periastron (Degrees) | 90 (fixed) | ||
| Equilibrium temperaturea (K) | |||
| Planet-to-star radius ratio | |||
| Semi-major axis in stellar radii | |||
| Transit depth in (ppt) | |||
| Transit depth in (ppt) | |||
| Transit depth in (ppt) | |||
| Transit depth in (ppt) | |||
| Transit depth in TESS (ppt) | |||
| Ingress/egress transit duration (days) | |||
| Total transit duration (days) | |||
| Transit impact parameter | |||
| Incident Flux () | |||
| Separation at mid-transit | |||
| Predicted Parameters: | |||
| Planet massb () | |||
| RV semi-amplitude (m/s) | |||
| TSM | Transmission spectroscopy metric | 60.0 | 50.0 |
| Planet-to-star mass ratio () | |||
| Planet mean density (cgs) | |||
aThe equilibrium temperature is estimated supposing null albedo and an 100% efficient heat recirculation. bThe planetary mass is estimated using the Chen & Kipping (2017) mass–radius relation. The radial velocity semi-amplitude () is predicted using the estimated mass.
| Parameter | Description | Circular | Eccentric |
| Stellar Parameters: | |||
| Mass () | |||
| Radius () | |||
| Luminosity () | |||
| Density (cgs) | |||
| Surface gravity (cgs) | |||
| Effective Temperature (K) | |||
| Planetary Parameters: | |||
| Orbital period (days) | |||
| Planet radius () | |||
| Transit time () | |||
| Semi-major axis (AU) | |||
| Inclination (Degrees) | |||
| Eccentricity | 0 (Fixed) | ||
| Tidal circularization timescale (Gyr) | |||
| Argument of Periastron (Degrees) | 90 (fixed) | ||
| Equilibrium temperaturea (K) | |||
| Planet-to-star radius ratio | |||
| Semi-major axis in stellar radii | |||
| Transit depth in (ppt) | |||
| Transit depth in (ppt) | |||
| Transit depth in (ppt) | |||
| Transit depth in TESS (ppt) | |||
| Ingress/egress transit duration (days) | |||
| Total transit duration (days) | |||
| Transit impact parameter | |||
| Incident Flux () | |||
| Separation at mid-transit | |||
| Predicted Parameters: | |||
| Planet massb () | |||
| RV semi-amplitude (m/s) | |||
| TSM | Transmission spectroscopy metric | 38.0 | 40.0 |
| Planet-to-star mass ratio () | |||
| Planet mean density (cgs) | |||
aThe equilibrium temperature is estimated supposing an null-albedo and a 100% efficient heat recirculation. bThe planetary mass is estimated using the Chen & Kipping (2017) mass–radius relation. The radial velocity semi-amplitude () is predicted using the estimated mass.

### v3

@@
We present the validation of two TESS super-Earth candidates transiting the mid-M dwarfs TOI-6002 and TOI-5713 every 10.90 and 10.44 days, respectively. The first star (TOI-6002) is located pc away, with a radius of , a mass of , and an effective temperature of K. The second star (TOI-5713) is located pc away, with a radius of , a mass of , and an effective temperature of K. We validated the planets using TESS data, ground-based multi-wavelength photometry from many ground-based facilities, as well as high-resolution AO observations from Keck/NIRC2. TOI-6002 b has a radius of and receives . TOI-5713 b has a radius of and receives . Both planets are located near the radius valley and near the inner edge of the habitable zone of their host stars, which makes them intriguing targets for future studies to understand the formation and evolution of small planets around M-dwarf stars.
@@
The abundance and the distribution of planets with radii below 4 for orbits shorter than 100 days allows the identification of specific features that can link their abundance to their formation and evolution like the radius valley (see e.g. Fulton et al., 2017; Cloutier & Menou, 2020) and the density valley (Luque & Pallé, 2022), based on their radii and densities distributions. The radius valley features a lack in planets ranging from 1.5 to 2 . In this regard, some studies argue that rocky super-Earths are a byproduct of atmospheric escape experienced by gaseous sub-Neptunes driven by the energy radiated by the host stars (see e.g. Lopez & Fortney, 2013; Owen & Wu, 2013; Jin et al., 2014; Chen & Rogers, 2016) and (or) the energy released by their cooling planetary cores (see e.g. Owen & Wu, 2013; Ginzburg et al., 2016, 2018). Other studies argue that some planets are formed as rocky super-Earths in a gas-depleted (see e.g. Lopez & Rice, 2018; Cloutier & Menou, 2020; Cherubim et al., 2023) or gas-poor environment with thin gaseous envelopes (see e.g. Lee & Connors, 2021; Cherubim et al., 2023). Additional theoretical studies predict that the sub-population of sub-Neptunes consists typically of water-rich worlds formed beyond the water ice-line and migrate toward their host stars (se e.g. Venturini et al., 2020; Burn et al., 2024; Venturini et al., 2024). The density valley features three peaks in the density distribution defining rocky super-Earths, water-rich worlds, and puffy sub-Neptunes. Luque & Pallé (2022) suggest that the pebble accretion is the main mechanism forming these planets with super-Earths formed within the ice line and water-rich world formed outside it (se e.g. Venturini et al., 2020; Burn et al., 2024; Venturini et al., 2024). For now the mechanisms forming super-Earths remain unclear, making observations of planets in this mass-radius regime a critical part to understand it.
TESS is increasing the number of small exoplanets amenable to mass and atmospheric characterization (Ricker et al., 2015; Sullivan et al., 2015). Detailed studies of such exoplanets will allow insights into the origin of the radius and density valleys in the near future. Such analyses should shed further light on the mechanisms dominating the formation and evolution of super-Earths. In addition, the analyses of exoplanets with incident irradiation between Earth and Venus will allow key insights into the limits of the HZ (e.g. Kaltenegger et al., 2023a).
@@
The paper is structured as follows. Sect. 2 introduces the data from TESS and all ground-based observations. Stellar characterization, validation of the planetary nature of the two TESS candidates, and global data analyses are covered in Sect. 3. Results, potential for mass measurements, and atmospheric characterization are discussed for both planets in Sect. 4. Sect. 5 summarizes our findings.
@@ 2 Observations
Aperture size column presents the optimum uncontaminated apertures used for extracting the light curves. Detrending parameters were chosen on the basis of maximizing the log-likelihood of each transit.
@@ 2.1 TESS photometry
The TESS time-series observations were processed by the TESS Science Processing Operations Center (SPOC) pipeline (Jenkins et al., 2016, 2020). SPOC calibrates the image data, carries out quality control, labels bad data, and performs simple aperture photometry (SAP) for each star in the field of view. Then, it corrects for instrumental systematic offsets using a Presearch Data Conditioning module (PDC: Smith et al., 2012; Stumpe et al., 2012, 2014) and produces pre-search data conditioned light curves. Appendix 12 shows the target pixel files (TPFs) and the SAP apertures used for photometry, along with the overlaid locations of nearby Gaia DR2 (Gaia Collaboration et al., 2018) sources.
@@
For our analyses, we downloaded the 2-minute PDC-SAP light curves from the Mikulski Archive for Space Telescopes (MAST111https://archive.stsci.edu/). Using the ”QUALITY” flag, we identified all the bad-quality data points that were removed from the data files. The year two data of TESS were subject to sky background overestimation in the SPOC pipeline for dim and (or) crowded targets, affecting the transit depth and, thus, the planet radius. TOI-6002 is a crowded-field target and was subjected to this effect in sector 14. We corrected the PDC-SAP light curves of sector 14 using equation 2 described in section 4.2 of the data release notes of sector 27222https://tasoc.dk/docs/release_notes/tess_sector_27_drn38_v02.pdf (see section 4.2, equation 2). This effect was negligible for TOI-5713 due to the low level of crowding. The light curve of TOI-5713 showed some clear flare-like signals, which we removed using an upper 3 clipping. We used wotn (Hippke et al., 2019) to detrend the light curves for stellar variability, utilizing a bi-weight filter with a window 2.5 times longer than each planets’ transit duration. The obtained lightcurves are presented in Appendix 13, with the locations of the transit signals highlighted with different colors.
@@ 2.2.1 LCOGT-1m0 photometry
We used the 40964096 SINISTRO CCD camera on the 1.0-m robotic telescope of the Las Cumbres Observatory Global Telescope (LCOGT; Brown et al. 2013) network node at Teide Observatory (Tenerife, Spain) to observe two full transits of TOI-6002 b on UT 2023 May 13 and 24. Table 1 provides the observation parameters. We used the standard LCOGT BANZAI pipeline (McCully et al., 2018) to calibrate the science images, and we used AstroImageJ (AIJ: Collins et al., 2017) for the photometric extraction.
@@ 2.2.4 MuSCAT2 photometry
For TOI-5713 b, we observed one full transit on UTC July 8, 2023 (see Table 1), using the multi-color imager (, , , and bands) MuSCAT2 (Narita et al., 2019), mounted on the 1.52-m Telescopio Carlos Sánchez (TCS) at Teide Observatory, Spain. We used the band images for guiding, as the star is faint in the blue optical ( mag). We calibrated and reduced the raw data using the code and procedure of Parviainen et al. (2019). Due to the short exposure time and the faintness of the target, the MuSCAT2 band photometry presented a much larger scatter than in the other bands, and it was not used in the analysis presented below.
@@ 2.3.1 IRTF/SpeX
We used the SpeX spectrograph (Rayner et al., 2003) on the 3.2-m NASA Infrared Telescope Facility (IRTF) to collect a near-infrared spectrum of TOI-5713 on 2023 May 16 (UT) and TOI-6002 on 2023 Oct 31 (UT). On both nights, conditions were clear with seeing of 06–07. We used the short-wavelength cross-dispersed (SXD) mode, and aligned the slit to the parallactic angle. The resulting spectra span 0.80–2.42 m and have a resolving power of . We collected six integrations of 139.9 s on TOI-5713 and 299.8 s on TOI-6002, using an ABBA nod pattern. Immediately after the science observations, we gathered a standard set of flat-field and arc-lamp calibrations and observed A0 V standard stars. We reduced the data with Spextool v4.1 (Cushing et al., 2004), following the standard instructions from the Spextool User’s Manual333Available at http://irtfweb.ifa.hawaii.edu/~spex/observer/.. The final spectrum of TOI-5713 has a median per-pixel S/N of 92 with peak values in the , , and bands of of 116, 131, and 121, respectively. For TOI-6002, the final spectrum has a median S/N per pixel of 156 with peak values of 195, 219, and 205 in the , , and bands, respectively.
@@ 2.4 High-resolution imaging
We assessed the sensitivities of the final combined AO image by injecting simulated sources azimuthally around the primary target, following the approach of Furlan et al. (2017). Fig. 1 shows the resulting sensitivity curves. We do not detect close-in () stellar companions for either star. The Keck and Palomar data have sensitivities of mag and mag at 0.5″, respectively. This corresponds to a spatial limit au for TOI-6002 and au for TOI-5713.
@@ 3.2 Discarding possible false-positive scenarios
The second astrophysical false positive case is that of bound eclipsing binaries. Many findings argue against this scenario for both stars: (1) the spectral energy distribution is fit well by single-star models (see Fig. 4); (2) the spectra collected by SpeX SXD and Shane/Kast, for both stars, show no secondary spectrum (see Sect. 2), and Fig. 2 and 3; (3) the achromatic nature of the transits, where their depths in all bands agree within for each planet (see Table 5); (4) the stellar densities determined from the transits alone are in good agreement with the those determined from the the basic parameters of each star (see Table 3). Considering these arguments, we discard these false-positive scenarios.
@@ 3.6 Global modeling
In the priors file, we set starting values of , and . We set lower and upper limits priors on the orbital period ( d) and the transit epoch () from the values determined by SPOC and reported in ExoFOP. For the LD parameters, we disabled the priors based on the Claret tables (Claret, 2017; Claret & Bloemen, 2011) by setting the NOCLARET flag given the low mass of our stars where these tables are not reliable (see Eastman et al., 2019, and references therin). Instead, we fitted for quadratic LD parameters assuming Gaussian priors of variance 0.1, computed with the PyLDTk code (Parviainen, 2015) for each passband (see Table 4). Regarding the TESS data, we used the detrended PDC-SAP light curves (see Sect. 2.1) which are corrected for the contribution of neighboring stars and subsequently corrected for sky background bias for TOI-6002 in sector 14 (see Sect. 2.1), but we still fit for a possible dilution by applying a Gaussian prior of 010% of the contamination ratio reported in ExoFOP. This is to propagate any uncertainty in the correction. Also, Gaussian priors were applied on the metallicities from IRTF/SpeX (see Table 3). For each planet, we fitted for circular and eccentric orbits to check for orbital eccentricity evidences using only photometric data. The two fits were run until convergence. We found the difference in the loglikelihood is for both TOI-6002 b and TOI-5713 b, which suggests that eccentric orbits are marginally favored for both planets. Table 5 presents the fits’ results for both the circular and eccentric cases. Fig. 8 shows the detrended and modeled gound-based light curves for both TOI-6002 b and TOI-5713 b, and Fig. 9 shows the TESS phase-folded light curves.
@@
| Parameter | Description | TOI-6002 b | TOI-5713 b | ||
| Circular | Eccentric | Circular | Eccentric | ||
| Stellar Parameters: | |||||
| Mass () | |||||
| Radius () | |||||
| Luminosity () | |||||
| Density (cgs) | |||||
| Surface gravity (cgs) | |||||
| Effective Temperature (K) | |||||
| Planetary Parameters: | |||||
| Orbital period (days) | |||||
| Planet radius () | |||||
| Transit time () | |||||
| Semi-major axis (AU) | |||||
| Inclination (Degrees) | |||||
| Eccentricity | 0 (Fixed) | 0 (Fixed) | |||
| Tidal circularization timescale (Gyr) | |||||
| Argument of Periastron (Degrees) | 90 (fixed) | 90 (fixed) | |||
| Equilibrium temperaturea (K) | |||||
| Planet-to-star radius ratio | |||||
| Semi-major axis in stellar radii | |||||
| Transit depth in (ppt) | … | … | |||
| Transit depth in (ppt) | |||||
| Transit depth in (ppt) | |||||
| Transit depth in (ppt) | |||||
| Transit depth in TESS (ppt) | |||||
| Ingress/egress transit duration (days) | |||||
| Total transit duration (days) | |||||
| Transit impact parameter | |||||
| Incident Flux () | |||||
| Separation at mid-transit | |||||
| Predicted Parameters: | |||||
| Planet massb () | |||||
| RV semi-amplitude (m/s) | |||||
| TSM | Transmission spectroscopy metric | 60.0 | 50.0 | 38.0 | 40.0 |
| Planet-to-star mass ratio () | |||||
| Planet mean density (cgs) | |||||
aThe equilibrium temperature is estimated supposing a null-albedo and a 100% efficient heat recirculation. bThe planetary mass is estimated using the Chen & Kipping (2017) mass–radius relation. The radial velocity semi-amplitude () is predicted using the estimated mass.